OpenStudy (anonymous):

Write the equation for the line with the given properties passing through (-5,7) and (3,-4)

OpenStudy (anonymous):

for this you should use the point-slope equation

OpenStudy (anonymous):

get slope first

OpenStudy (anonymous):

(y-y1)=m(x-x1) where m is the slope and y and x are coordinates of a point on the line

OpenStudy (anonymous):

to find the slope use\[\frac{ y _{1}-y _{2} }{ x _{1}-x _{2} }\]

OpenStudy (anonymous):

I'm kind of just checking my answer...should i come out to : y=-11/8x+89/8

OpenStudy (anonymous):

just plug in the points

OpenStudy (anonymous):

Can someone please tell me if I got the correct answer? Thanks.

OpenStudy (anonymous):

I got y = -11/8x+1/8